# Adversarial Inception for Bounded Backdoor Poisoning in Deep Reinforcement Learning

### v1

@@ 3.3 Threat Model
In this work, in line with the existing online backdoor literature, we assume that the adversary is able to infiltrate the system on which the agent is being trained such that they can run processes on the machine and manipulate RAM values. While these are strong assumptions, they are unfortunately realistic as there have been at least 5,175 incidents of system intrusions in 2024 alone as reported by the most recent Verizon DIBR report (Hylender et al., 2024).
We will also be leveraging the “outer-loop” threat model used by the SleeperNets attack since it provides increased versatility for the adversary over the prior “inner-loop” attacks like TrojDRL. Under this threat model, our adversary can observe episodes $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$ of size $\mu$ completed by the agent in $M$. The adversary can then alter states $s_{t}$, actions $a_{t}$, and rewards $r_{t}$ stored in the trajectory before the agent uses them in their policy optimization. We note that here, unlike the action manipulation implemented in TrojDRl, our adversary changes actions after the episode has finished meaning these new actions will not actually occur in the environment. This makes our attack’s action manipulation much stealthier as the relationship between states and actions would need to be well defined and analyzed for any manipulation to be detected.
The adversary is also constrained by a poisoning rate parameter $\beta$ which bounds the proportion of states $s$ the adversary can apply the trigger to. These constraints are standard throughout poisoning literature in machine learning (Jagielski et al., 2021). In DRL $\beta$ acts similar to a hyper parameter for the adversary. At lower values the adversary poisons less states, allowing the agent to more easily optimize the benign MDP, but potentially decreasing the attack’s success rate. At higher values the adversary poisons more states, likely leading to an increase in attack success rate, but potentially decreasing the agent’s performance in the benign task or making the attack easier to detect.
@@ 4 Theoretical Analysis
In this section we formally propose and define a new class of backdoor poisoning attacks we refer to as “Adversarial Inception” attacks. We first explore prior attempts at action manipulation seen in attacks like TrojDRL, and show why they are ineffective at increasing the adversary’s attack success rate. We then formally define our proposed inception attack framework and present strong theoretical guarantees for attack success and attack stealth while satisfying our reward constraints.
@@ 4.1 Prior Action Manipulation is Ineffective
In prior works, such as TrojDRL, attempts were made to improve attack performance via action manipulation which occasionally forces the agent to take the target action in poisoned states at training time. We call this approach “forced action manipulation”. In this section we show how these approaches are inconsequential towards maximizing attack success since they merely influence the agent’s exploration. We model forced action manipulation as an adversarial policy $\pi^{+}$ which alters the agent’s true policy $\pi$ such that they are forced to take action $a^{+}$ with probability $\rho$:
@@
for some poisoned state $s_{p}\in S_{p}$. We now show how this does not provide any guarantees of attack success. Let’s again consider the MDP defined in Figure 3 under a backdoor attack using bounded reward poisoning, as defined in subsection 3.2, and forced action manipulation, as defined in Equation 6. We can then evaluate the value of each action within the adversarial state $\delta(\text{Start})$ in Table 3.
Q-Values With Action Manipulation Without Action Manipulation
@@
Table 3: Q values for an arbitrary policy $\pi$ in our example MDP (Figure 3) under a backdoor attack with and without forced action manipulation. Here $a^{+}$ is not optimal even with action manipulation.
With or without forced actions the value of $a^{+}$ in $\delta(\text{Start})$ is $1$ as the agent receives an immediate reward of $+1$ then terminates. Similarly, both with or without forced actions, when the agent takes action $a$ in $\delta(\text{Start})$ they receives an immediate reward of 00 followed by a reward of $+1$ on each subsequent time step, for an overall return of $\frac{\gamma}{1-\gamma}$. Neither of these terms depend on $\pi^{+}$, thus the value of $\pi$ in the poisoned state $\delta(\text{Start})$ does not change with the inclusion of $\pi^{+}$. Note that, since the values of the “Terminate” and “Propser” states are fixed with respect to any $\pi$, this result holds for both on and off policy learning. Thus, $a^{+}$ will not be the optimal action for any $\gamma>\frac{1}{2}$ and attack success is not attained. In more complex MDPs $\pi^{+}$ will have some influence on the value of the agent’s next state, however the key observation here is that forced action manipulation has no direct impact on the optimality of the target action $a^{+}$ in poisoned states.
4.2 Modeling Adversarial Inception
Before formulating Adversarial Inception there are a two assumptions, shared by SleeperNets, we make to ensure our theoretical results are maintained. First, we assume that $S_{p}$ and $S$ are disjoint. This ensures that any reward or action manipulation which occurs in a poisoned state does not conflict with an equivalent benign state. This assumption is reasonable in practice as the trigger can be designed to be distinct relative to the agent’s usual states. We additionally assume that the trigger function $\delta$ forms a bijection between $S$ and $S_{p}$, allowing us to invert the function to retrieve the benign state $s=\delta^{-1}(s_{p})$. In practice this assumption is unnecessary as the adversary observes $s$ first and then applies the trigger $s_{p}=\delta(s)$, thus they never need to compute an inversion of $\delta$.
With these assumptions in mind, we now formulate adversarial inception. In subsection 3.2 we explored how backdoor attacks using unbounded reward poisoning are susceptible to detection and how reward clipping negates their theoretical guarantees. In subsection 4.1 we further showed how “forced action manipulation” techniques are ineffective at increasing attack success rates. Thus, we propose our new adversarial inception framework which overcomes these shortcomings of prior attacks under bounded reward poisoning restrictions. To achieve this we design an adversarial MDP $M^{\prime}=(S\cup S_{P},A,R^{\prime},T^{\prime},\gamma)$ with a novel transition function $T^{\prime}$. Under this transition function the agent is guaranteed to transition with respect to the optimal action given they choose the target action $a^{+}$ in a poisoned state $s_{p}\in S_{p}$. We enforce this through the function $\phi$ below
@@ 4.2 Modeling Adversarial Inception
$\phi(s_{p},a,\pi)=\left\{\begin{array}[]{c|c}\operatorname*{arg\,max}_{a^{\prime}}[Q^{M^{\prime}}_{\pi}(\delta^{-1}(s_{p}),a^{\prime})]&\text{if }a=a^{+}\\ a&\text{otherwise}\end{array}\right.$ (8)
The goal of $\phi$ is to calculate and return the optimal action in state $\delta^{-1}(s_{p})$ given $\pi$ if the agent chooses the target action $a^{+}$. In practice the adversary doesn’t have direct access to $Q\pi^{M^{\prime}}$, so approximation is necessary. This will be explored in more detail in Section 5. $\phi$ is then incorporated into the transition function $T^{\prime}$ defined below where $\Pi$ is the set of all valid policies in $M^{\prime}$:
@@
$T^{\prime}(s,a,s^{\prime},\pi)\doteq\left\{\begin{array}[]{c|c}(1-\beta)\cdot T(s,a,s^{\prime})&\textit{if }s\in S,\;s^{\prime}\in S\\ \beta\cdot T(s,a,\delta^{-1}(s^{\prime}))&\textit{if }s\in S,\;s^{\prime}\in S_{p}\\ \beta\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;\delta^{-1}(s^{\prime}))&\textit{if }s\in S_{p},\;s^{\prime}\in S_{p}\\ (1-\beta)\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;s^{\prime})&\textit{if }s\in S_{p},\;s^{\prime}\in S\\ \end{array}\right.$ (10)
Thus, given any $s\in S_{p}\cup S$ the MDP transitions to a poisoned state $s^{\prime}\in S_{p}$ with probability $\beta$ and otherwise transitions to a benign state. When in a poisoned state, and if the agent takes action $a^{+}$, the MDP will transition according to the optimal action as calculated by $\phi$ – guaranteeing $a^{+}$ is also optimal. Furthermore, all poisoned states $s,s^{\prime}\in S_{p}$ are inverted back to their benign form $\delta^{-1}(s)$. This maintains the underlying transition dynamics of the environment with respect to states and actions, allowing us to maintain optimality between policies in $M$ and $M^{\prime}$.
The next key piece of adversarial inception is our adversarial reward function $R^{\prime}$ whose goal is to give the agent extra rewards or penalties when they do or don’t choose $a^{+}$ in poisoned states respectively. However, as we discussed in Section 3, the adversary’s reward function must be bounded by the limits of the benign reward function $R$. To account for this we define the function $\tau$ to give the agent as much reward or penalty as possible while remaining within these bounds:
@@
$\tau(a,r,\hat{r})=\left\{\begin{array}[]{cc}\min[r+\frac{\hat{r}-L}{\gamma},U]&\text{if }a=a^{+}\\ \max[r-\frac{U-\hat{r}}{\gamma},L]&\text{otherwise}\end{array}\right.$ (12)
where $\hat{r}$ is the reward received on the previous time step. Similar to the unbounded reward poisoning approach of SleeperNets, we want to cancel out this $\tau$ term in the value of the prior benign state $s\in S$. This is to ensure that values in benign states aren’t biased by our reward poisoning approach, allowing us to maintain the optimal policies of $M$. Prior benign states will receive $\tau$ discounted by a factor of $\gamma$, hence the division by $\gamma$ present here. Thus we define $R^{\prime}$ as
@@
$R^{\prime}(s,a,s^{\prime},\hat{r})\doteq\left\{\begin{array}[]{c|c}R(s,a,s^{\prime})&\textit{if }s\in S,\;s^{\prime}\in S\\ R(s,a,\delta^{-1}(s^{\prime}))+\mathbb{E}_{a,r\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]&\textit{if }s\in S,s^{\prime}\in S_{p}\\ \tau(a,R(\delta^{-1}(s),\phi(s,a,\pi),s^{\prime}),\hat{r})&\textit{if }s\in S_{p}\end{array}\right.$ (14)
where $\mathbb{E}_{r,a\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]$ is used as our aforementioned bias correction. $R^{\prime}$ is designed to work in tandem with $T^{\prime}$ to guarantee attack success – when taking action $a^{+}$ in some poisoned state $s_{p}\in S_{p}$ the MDP will not only transition according to the optimal action as calculated by $\phi$, but the agent will also receive increased immediate reward as specified by $\tau$. In Table 4 we show how this attack formulation overcomes the weaknesses of bounded reward poisoning and forced action manipulation. Here we once again return to our example MDP from Figure 3, except now, under inception poisoning, $a^{+}$ is finally the optimal action in the poisoned state $\delta(\text{start})$ for any $\gamma$.
Q-Values With Inception Without Inception
@@
When the agent takes action $a^{+}$ under adversarial inception, and according to $\phi$, the agent will actually transition with respect to action $a$, receiving a future return of $\frac{\gamma}{1-\gamma}$. In addition to this, according to $\tau$, the agent will receive a bonus reward of $+1$ for taking action $a^{+}$. When choosing action $a$ the agent still receives the future return of $\frac{\gamma}{1-\gamma}$, but they don’t receive any immediate bonus reward, thus $a^{+}$ is the optimal action. With no adversarial inception the attack can no longer transition the agent with respect to $a$ upon choosing $a^{+}$, thus they can only give the agent an immediate reward of $+1$, making $a$ the optimal action. We claim that this result generalizes across MDPs, allowing us to prove that $M^{\prime}$ not only maximizes attack success but also attack stealth, all while providing rewards that stay within the bounds of $R$. In the next section we will formalize these claims, all of which have proofs provided in the appendix.
@@ 4.3 Theoretical Guarantees of Adversarial Inception
Here we present the theoretical guarantees afforded to us by adversarial inception, with Theorem 1 and Theorem 2 relating to attack success and attack stealth respectively. The outcome of Theorem 1 is fairly intuitive based upon our prior explanations of $T^{\prime}$ and $R^{\prime}$ – if the agent is guaranteed an optimal outcome when choosing $a^{+}$ in poisoned states, then $a^{+}$ is always optimal.
Theorem 2, on the other hand, isn’t as obvious. When performing action manipulation according to $\phi$ it’s unclear how this will impact the dynamics of the MDP and thus the optimal policy. Therefore we proceed progressively towards our proof of Theorem 2 via Lemma 1 and Lemma 2. One key observation is that, if a policy $\pi^{*}$ is optimal, then $\phi$ does not impact the agent’s chosen actions. Thus Lemma 1 is the result. Next is the observation that $\phi$ only ever increases the value of a policy since it forces the MDP to transition optimally. Thus Lemma 2 follows.
@@
Through these lemmas we are given a direct relationship between the value of a policy $\pi$ in the benign MDP $M$ and the adversarial MDP $M^{\prime}$. With this we can prove Theorem 2. Therefore, given Theorem 1 and Theorem 2 we know an optimal policy in $M^{\prime}$ solves both our objectives of attack success and attack stealth while satisfying our reward constraints. Therefore, since DRL algorithms are designed to converge towards an optimal policy, we know that adversarial MDPs, $M^{\prime}$, designed according to adversarial inception will solve both attack success and stealth. Formally derived proofs for all these results are given in the appendix.

### v3

@@ 3.3 Threat Model
In this work we consider the outer-loop threat model defined by SleeperNets, which provides the attacker global visibility into a history of observations before modifying states, actions and rewards. The outer-loop threat model, while assuming the same level of adversarial access to the training process as the inner-loop threat model (Kiourti et al., 2019), offers increased versatility (Rathbun et al., 2024). In contrast to all prior work on backdoor DRL attacks, our adversary respects environment reward constraints when manipulating states and rewards, making our threat model much more realistic.
Under our threat model, the adversary observes episodes $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$ of size $\mu$ generated by the agent in $M$ during training. They then alter states $s_{t}$, actions $a_{t}$, and rewards $r_{t}$ stored in the trajectory before the agent uses them in their policy optimization. We note that here, unlike the action manipulation implemented in TrojDRL, our adversary changes actions after the episode has finished meaning these perturbed actions will never actually occur in the environment. The adversary is also constrained by a poisoning rate parameter $\beta$ which bounds the proportion of total training time steps in which the adversary can apply the trigger to the agent’s current state. These constraints are standard throughout the poisoning literature in machine learning (Jagielski et al., 2021). In DRL $\beta$ acts similar to a hyper parameter for the adversary. At lower values the adversary poisons fewer time steps, allowing the agent to more easily optimize the benign MDP, but potentially decreasing the attack’s success rate. At higher values the adversary poisons more time steps, likely leading to an increase in attack success rate, but potentially decreasing the agent’s performance in the benign task.
@@ 4 Theoretical Analysis
Here we first prove that prior reward and action manipulation techniques are insufficient for attack success. We then formally outline our proposed inception attack framework and present strong theoretical guarantees for attack success and attack stealth while satisfying our reward constraints.
@@ 4.1 Prior Action Manipulation is Ineffective
In Section 3.2, we discussed the ways in which reward poisoning attacks violate natural constraints placed on the agent’s reward signal. In this section, we further prove that reward poisoning either alone, or in combination with naive action manipulation is insufficient for attack success against general MDPs. Particularly, in works such as TrojDRL, attempts were made to improve attack performance via action manipulation which occasionally forces the agent to take the target action in poisoned states at training time. We call this approach “forced action manipulation” and model it as an adversarial policy $\pi^{+}$ which alters the agent’s true policy $\pi$ so they are forced to take action $a^{+}$ with probability $\rho$ in poisoned states, otherwise leaving the policy unmodified:
@@
for some poisoned state $s_{p}\in S_{p}$ where $\mathbbm{1}$ is the indicator function. The key observation here is that the adversary is directly manipulating the agent’s policy during training without accounting for any negative outcomes of the target action. Subsequently, forced action manipulation makes no contribution towards increasing the expected return of the target action in any poisoned state $\delta(s)$.
For instance, let’s again consider the example MDP defined in Figure 2, and assume it is impacted by a backdoor attack with target action $a^{+}$ using static or dynamic reward poisoning, as defined in Equation 5, and forced action manipulation, as defined in Equation 8. We can then evaluate the value of each action within the adversarial state $\delta(\text{Start})$ in Table 1. Since rewards in this MDP are constrained to $[0,1]$, both static and dynamic reward poisoning will give the agent a reward of $1$ if they take action $a^{+}$ or 00 otherwise. Thus, we need not analyze them separately. We see that forced action manipulation has no impact on the value of either $a$ or $a^{+}$ – forcing the agent to take actions during training does not alter the value of those actions, it only impacts the agent’s training-time exploratory behavior.
Q-Values Manipulation No Manipulation
@@
Table 1: Q-values for an arbitrary policy $\pi$ in our example MDP (Figure 2) under a constrained backdoor attack with and without forced action manipulation. Here $a^{+}$ is suboptimal with and without action manipulation for any $\gamma\geq\frac{1}{2}$.
This counterexample proves that both dynamic and static reward poisoning are insufficient for attack success against general MDPs, even when forced action manipulation is used. Note that, since the values of the “Terminate” and “Prosper” states are fixed with respect to any $\pi$ and the Q-value provides the first action taken, this result holds for both on and off policy learning.
4.2 Adversarial Inception Framework
In this section, we formalize our new adversarial inception framework which overcomes the shortcomings of prior attacks when subject to reward poisoning constraints. To achieve this, we design an adversarial MDP with a novel transition function $T^{\prime}$ and reward function $R^{\prime}$. Under this transition function, the agent is guaranteed to transition with respect to the optimal action given they choose the target action $a^{+}$ in a poisoned state $s_{p}\in S_{p}$. We enforce this through the function $\phi$ below
@@ 4.2 Adversarial Inception Framework
$\phi(s_{p},a,\pi)=\left\{\begin{array}[]{c|l}\operatorname*{arg\,max}_{a^{\prime}}[Q^{M^{\prime}}_{\pi}(\delta^{-1}(s_{p}),a^{\prime})]&a=a^{+}\\ a&\text{otherwise}\end{array}\right.$
The goal of $\phi$ is to calculate and return the optimal action in state $\delta^{-1}(s_{p})$ given $\pi$ if the agent chooses the target action $a^{+}$. In practice, the adversary doesn’t have direct access to $Q\pi^{M^{\prime}}$, so approximation is necessary. This will be explored in more detail in Section 5. Further note that for our theoretical results we assume $S$ and $S_{p}$ are disjoint sets and $\delta$ is a bijection between them. The adversary never needs to compute $\delta^{-1}(s_{p})$ in practice, however, as they can directly observe $s$ before the trigger is applied. $\phi$ is then incorporated into the transition function $T^{\prime}$ defined below where $\Pi$ is the set of all valid policies in $M^{\prime}$:
@@
Adversarial Transition Function $T^{\prime}$
$s$ $s^{\prime}$ Transition Probability
$S$ $S$ $(1-\beta)\cdot T(s,a,s^{\prime})$
$S$ $S_{p}$ $\beta\cdot T(s,a,\delta^{-1}(s^{\prime}))$
$S_{p}$ $S$ $(1-\beta)\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;s^{\prime})$
$S_{p}$ $S_{p}$ $\beta\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;\delta^{-1}(s^{\prime}))$
Thus, given any $s\in S_{p}\cup S$ the MDP transitions to a poisoned state $s^{\prime}\in S_{p}$ with probability $\beta$ and otherwise transitions to a benign state. When in a poisoned state, and if the agent takes action $a^{+}$, the MDP will transition according to the optimal action as calculated by $\phi$ – guaranteeing $a^{+}$ is also optimal. Furthermore, all transition probabilities are computed with respect to the benign states $\delta^{-1}(s_{p})$ for $s_{p}\in S_{p}$. This maintains the underlying transition dynamics of the environment with respect to states and actions, retaining optimality between policies in $M$ and $M^{\prime}$.
The next key piece of adversarial inception is our adversarial reward function $R^{\prime}$ whose goal is to give the agent extra rewards or penalties when they do or don’t choose $a^{+}$ in poisoned states, respectively. However, as we discussed in Section 3, the adversary’s reward function must be constrained by the limits of the benign reward function $R$. To account for this we define the function $\tau$ to give the agent as much reward or penalty as possible while remaining within the bounds $[L,U]$ where $L=\inf(R)$, $U=\sup(R)$:
@@
$\tau(a,r,\hat{r})=\left\{\begin{array}[]{cc}\min[r+\frac{\hat{r}-L}{\gamma},U]&\text{if }a=a^{+}\\ \max[r-\frac{U-\hat{r}}{\gamma},L]&\text{otherwise}\end{array}\right.$
where $\hat{r}$ is the reward received on the previous time step. We want to cancel out this $\tau$ term in the value of the prior benign state $s\in S$. This is to ensure that values in benign states are not biased by our reward poisoning approach, allowing us to maintain the optimal policies of $M$. Returns in prior benign states will receive $\tau$ discounted by a factor of $\gamma$, hence the division by $\gamma$ present here. Thus we define $R^{\prime}$ as
@@
Adversarial Reward Function $R^{\prime}$
$s$ $s^{\prime}$ Reward
$S$ $S$ $R(s,a,s^{\prime})$
$S$ $S_{p}$ $R(s,a,\delta^{-1}(s^{\prime}))+\gamma\mathbb{E}_{a,r\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]$
$S_{p}$ $S\cup S_{p}$ $\tau(a,R(\delta^{-1}(s),\phi(s,a,\pi),s^{\prime}),\hat{r})$
where $\gamma\mathbb{E}_{r,a\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]$ is used as our aforementioned bias correction. $R^{\prime}$ is designed to work in tandem with $T^{\prime}$ to guarantee attack success – when taking action $a^{+}$ in some poisoned state $s_{p}\in S_{p}$ the MDP will not only transition according to the optimal action as calculated by $\phi$, but the agent will also receive an increased immediate reward as specified by $\tau$. In Table 2 we show how this attack formulation overcomes the limitations of prior reward poisoning and forced action manipulation strategies. Here, we once again return to our example MDP from Figure 2, except now, under inception poisoning, $a^{+}$ is finally the optimal action in the poisoned state $\delta(\text{start})$ for any $\gamma$.
Q-Values Inception No Inception
@@
When the agent takes action $a^{+}$ under adversarial inception, and according to $\phi$, the agent will instead transition with respect to action $a$, receiving a future return of $\frac{\gamma}{1-\gamma}$. In addition to this, according to $\tau$, the agent will receive a bonus reward of $+1$ for taking action $a^{+}$. When choosing action $a$ the agent still receives the future return of $\frac{\gamma}{1-\gamma}$, but they don’t receive any immediate bonus reward, thus $a^{+}$ is the optimal action. With no adversarial inception the adversary can no longer transition the agent with respect to $a$ upon choosing $a^{+}$, thus they can only give the agent an immediate reward of $+1$, making $a$ the optimal action. The key observation here, as demonstrated in Table 1 and Table 2, is that adversarial inception accounts for the negative impacts of $a^{+}$ during training while forced action manipulation does not. We claim that this result generalizes across MDPs, allowing us to prove that optimal policies in $M^{\prime}$ not only maximize attack success but also attack stealth, all while receiving rewards that stay within the bounds of $R$. We formalize these claims in the next section.
@@ 4.3 Theoretical Guarantees of Adversarial Inception
Here, we present the theoretical guarantees we have proven about adversarial inception, with Theorem 1 and Theorem 2 relating to attack success and attack stealth, respectively. The outcome of Theorem 1 is fairly intuitive based upon our prior explanations of $T^{\prime}$ and $R^{\prime}$ – if the agent is guaranteed an optimal outcome when choosing $a^{+}$ in poisoned states, then $a^{+}$ is always optimal.
Theorem 1 $\operatorname*{arg\,max}_{a}[Q^{M^{\prime}}_{\pi}(s_{p},a)]=a^{+}\;\forall s_{p}\in S_{p},\pi\in\Pi$. Thus, the optimal action of any policy in $M^{\prime}$ in any poisoned state $s_{p}$ is the targeted action $a^{+}$.
Theorem 2, on the other hand, is not as obvious. When performing action manipulation according to $\phi$ it is unclear how this will impact the dynamics of the MDP and thus the optimal policy. Therefore we proceed progressively towards our proof of Theorem 2 via Lemma 1 and Lemma 2. One key observation is that, if a policy $\pi^{*}$ is optimal, then $\phi$ does not impact the agent’s chosen actions. Thus Lemma 1 holds. Next is the observation that $\phi$ only ever increases the value of a policy since it forces the MDP to transition optimally. Thus Lemma 2 follows.
Lemma 1 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S\cup S_{p},\pi\in\Pi\Rightarrow V^{M^{\prime}}_{\pi^{*}}=V^{M}_{\pi^{*}}$ Therefore if $\pi^{*}$ is optimal then its value in $M^{\prime}$ is equal to its value in $M$.
@@
Through these lemmas we now have a direct relationship between the value of a policy in the benign MDP $M$ and the adversarial MDP $M^{\prime}$. With this we can prove Theorem 2.
@@
Therefore, given Theorem 1 and Theorem 2 we know an optimal policy in $M^{\prime}$ solves both our objectives of attack success and attack stealth while satisfying our reward constraints. Therefore, since DRL algorithms are designed to converge towards an optimal policy, we know that adversarial MDPs, $M^{\prime}$, designed according to adversarial inception will solve both attack success and stealth. Formally derived proofs for all these results are given in Appendix A.1.